OpenStudy (anonymous):

exact value of tan 75

OpenStudy (dumbcow):

\[\tan(a+b) = \frac{\tan a + \tan b}{1 - \tan a \tan b}\] 45+30 = 75

OpenStudy (solomonzelman):

or you can draw a diagram with 75 degree angle....